# TennisTV: Do Multimodal Large Language Models Understand Tennis Rallies?

### v1

@@ Abstract
Multimodal large language models (MLLMs) excel at general video understanding but struggle with fast, high-frequency sports like tennis, where rally clips are short yet information-dense. To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding. TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. It covers 8 tasks at rally and stroke levels and includes 2,500 human-verified questions. Evaluating 16 representative MLLMs, we provide the first systematic assessment of tennis video understanding. Results reveal substantial shortcomings and yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
@@ 1 Introduction
Fig. 1: Illustrations of a simple rally counting task easily handled by non-experts, but still challenging for the advanced commercial models.
In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos. TennisTV defines a stroke event as the atomic unit, spanning movement initiation, stroke action choice, directional execution, and the ball’s landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events. A total of 8 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements. We design an automated annotation pipeline. Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[5], extract additional athletes’ movement information, and generate questions. Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2,500 curated questions.
@@ 2.1 Sports Understanding
Sports understanding [6] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [7], athlete analysis [8], tactical planning [9], foul recognition [10]. Recent work on MLLMs [11, 12, 13] has shifted toward general frameworks spanning diverse sports tasks, while existing work[5] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally-level context and causal relations. Consequently, MLLM capabilities in fast-paced, high-frequency sports video remain largely unexplored. This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
@@ 3 TennisTV Benchmark
To rigorously evaluate the performance of models in tennis video analysis, we introduce Tennis Tour Video (TennisTV). The benchmark comprises 8 subtasks with 2500 questions, supporting multiple levels of video understanding. As shown in Fig 2, this section outlines benchmark overview and automatic annotation pipeline.
Table 1: Main results on TennisTV. AvgSE, AvgRS denotes the average metrics of tasks related to stroke event and rally sequence.
Method Frame Sampling AR MG HD HO AvgSE TP RC TI PW AvgRS Avg
Random - 25.0 20.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 31.3 27.9
@@
Video-LLaVA-7B 8 Frames 25.2 19.4 16.3 32.3 23.3 23.7 23.3 25.2 50.0 30.6 26.9
LLaVA-OV-7B 32 Frames 36.3 14.0 30.0 41.0 30.3 36.3 25.3 40.4 49.0 37.8 34.0
mPLUG-Owl3 128 Frames 42.8 16.1 33.8 46.0 34.7 35.7 18.3 39.3 50.0 35.8 35.3
Qwen2VL-7B 32 Frames 29.5 18.9 34.5 40.3 30.8 43.1 24.8 44.8 50.0 40.7 35.7
InternVid2.5 8B 128 Frames 49.5 16.7 17.8 43.0 31.8 42.2 19.0 42.6 50.3 38.5 35.1
Qwen2.5VL-3B 32 Frames 39.7 18.3 32.0 40.3 32.6 38.8 20.5 45.2 50.0 38.6 35.6
Qwen2.5VL-7B 32 Frames 48.6 28.3 34.5 45.7 39.3 41.9 22.3 45.6 48.0 39.5 39.4
MiMoVL-7B SFT 32 Frames 45.5 22.2 31.8 38.3 34.5 36.3 31.8 34.4 46.0 37.1 35.8
@@
Video-R1 32 Frames 45.9 22.8 28.8 42.0 34.9 43.7 28.3 43.7 52.0 41.9 38.4
VCR-7B 32 Frames 45.5 26.7 35.5 23.7 32.9 34.2 33.3 43.7 47.7 39.7 36.3
VideoRFT 32 Frames 46.8 21.7 35.3 45.0 37.2 35.1 25.5 42.6 45.7 37.2 37.2
VersaVid-R1 32 Frames 52.3 21.7 36.8 44.7 38.9 38.2 27.0 47.8 50.0 40.8 39.8
MiMoVL-7B RL 32 Frames 48.0 21.1 33.3 42.7 36.3 40.9 28.8 42.6 57.3 42.4 39.3
@@ 3.1 Benchmark Task Analysis
Stroke Event. The stroke-level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO). These tasks compel the model to detect subtle spatiotemporal cues at the frame level during rapid, interactive exchanges.
Rally Sequence. The rally-level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP) and Point Winner determination (PW). These tasks require a model to analyze the entire game process and perceive each stroke event, so as to grasp the overall trend of the rally and the players’ technical tendencies.
Overall, TennisTV offers a comprehensive and challenge benchmark that probes both basic perception (e.g., AR, HD) and higher-level tactical reasoning (e.g., PW, TP), providing a robust, diversified testbed for video understanding in fast-paced, high-frequency sports.
@@ 3.2 Dataset Annotation Pipeline
TennisTV is derived from F3Set[5], which provides manual annotations of individual shots from tennis tournament videos. Based on these annotations, we perform a three-stage data cleaning and annotation pipeline.
Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally-level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information. Next, we removed videos with an insufficient number of rallies from the original dataset. Moreover, because tennis tournaments are played on courts with diverse surfaces, the varying background colors can reduce viewing clarity. To address this issue, we employed Qwen2.5VL-7B[20] as a video quality judge model to filter out videos with poor visual quality.
@@
Question Generation We employed GPT-4o to aggregate the structured rally information and generated a comprehensive rally sequence description for each video. Using the rally descriptions and basic statistics, we generated 8 task questions using either GPT-4o or a predefined template. All questions were manually verified, resulting in a total of 2500 questions across 1298 videos.
@@ 4.1 Implementation Details
We evaluate 13 open-source MLLMs that can process multi-frame video inputs, including 8 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] . In addition, we assess GPT-4o, GPT-4.1 and Gemini2.5-Pro through their APIs. For the open-source models, we followed the input frame settings in VideoMME [24] . For the closed-source models, we adopted a video sampling rate of 1 FPS for evaluation due to budget consideration.
@@ 4.2 Evaluation Metric
We adopt two different metrics to evaluate the model performance on TennisTV. For PW task, which aims to identify which player won the rally, the model is provided with a statement describing the outcome of the rally. The model’s answer is considered correct only if it correctly classifies the statement as True or False. For other tasks, we formulate them as multiple-choice questions and evaluate whether the model’s selected option is correct.
@@ 4.3 Main Results
Results Table 3 reports results on TennisTV. Among closed-source models, GPT-4.1 ranks first overall at 45.0%, edging GPT-4o and Gemini 2.5 Pro. Open-source systems are competitive: the strongest non-thinking baseline, Qwen2.5-VL-7B, reaches 39.4%, and the thinking variant VersaVid-R1 nudges this to 39.8%. Reasoning brings targeted gains on the same 7B backbone (RC improves from 22.3% to 27.0% and TI from 45.6% to 47.8%) while averages remain close. Leadership is task-specific: Gemini 2.5 Pro leads MG (41.7%), GPT-4.1 leads TP (52.0%), and GPT-4o attains the highest AvgRS (48.1%). Despite these gains, overall accuracies remain modest and only moderately above the random baseline, indicating substantial headroom for progress.
Fig. 3: Correlation matrix between 8 tasks in TennisTV.
@@
1 FPS 37.5 39.2 38.3
16 Frames 38.1 39.8 39.0
32 Frames 39.3 39.5 39.4
64 Frames 38.7 38.5 38.6
Finding II — Optimal frame sampling density differs by task. As shown in Table 2, with Qwen2.5-VL-7B, overall performance peaks at 32 frames and further increasing the number of input frames yields diminishing or even negative returns, with the rally sequence metrics particularly affected. These patterns indicate that finer temporal resolution primarily benefits brief, high-frequency stroke cues, while rally-level reasoning saturates earlier and is more sensitive to redundancy.
Fig. 4: Visualization of performance on three TennisTV benchmark tasks, which require no prior knowledge of tennis, based on the same base model.
Finding III — Reasoning brings gains on knowledge-light tasks. To disentangle the effect of reasoning from tennis-specific priors, we analyze three knowledge-light tasks (MG, HD, RC) that are deliberately designed to minimize dependence on specialized domain knowledge (e.g., scoring rules). As shown in Fig. 4, three reasoning variants (VideoChat-R1, VideoRFT and VersaVid-R1) are derived from reinforcement learning post-training applied to the same Qwen2.5-VL-7B. With the backbone, tokenization scheme, and video interface held constant, the observed improvements over the non-reasoning baseline can be attributed to enhanced reasoning rather than architectural differences or domain-specific priors. In addition, we employ the chain-of-thought prompting method [25] to evaluate the original reasoning capability of the base model. Relative gains are most pronounced on RC and HD, where multi-step analysis and explicit temporal/trajectory reasoning are critical, and are steadier but smaller on MG. This pattern supports the view that explicit reasoning chiefly aids information aggregation, while offering more limited benefits on purely temporal-oriented tasks.
@@ 5 Takeaways
  • Enhancing the temporal understanding ability of video reasoning models. The experimental results have proved that reasoning can improve rally-level understanding, but it has not benefited from fine-grained temporal grounding task. This reveals that current video reasoning methods still lack sufficient capabilities for temporal perception.
@@ 6 Conclusion
In this paper, we introduce TennisTV, the first benchmark for tennis video understanding specifically designed to evaluate the capabilities of MLLMs. TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. Unified evaluations over diverse MLLMs show that reasoning and careful frame sampling help, yet models still struggle with the sport’s fast, information-dense dynamics. We expect TennisTV to spur advances in hierarchical temporal reasoning and robust sports video understanding.

### v2

@@ Abstract
Multimodal large language models (MLLMs) excel at general video understanding but struggle with fast, high-frequency sports like tennis, where rally clips are short yet information-dense. To systematically evaluate MLLMs in this challenging domain, we present TennisTV, the first and most comprehensive benchmark for tennis video understanding. TennisTV models each rally as a temporal-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. It covers 9 tasks from the stroke level to the rally level and includes 2943 human-verified questions. Evaluating 17 representative MLLMs, we provide the first systematic assessment of tennis video understanding. Results reveal substantial shortcomings and yield two key insights: (i) frame-sampling density should be tailored and balanced across tasks, and (ii) improving temporal grounding is essential for stronger reasoning.
@@ 1 Introduction
Fig. 1: Illustrations of a simple counting task easily handled by non-experts, but still challenging for the advanced commercial models.
In order to systematically evaluate the current MLLMs’ understanding of short, information-dense sports videos, as well as promote the community’s improvement of MLLMs’ understanding of tennis videos, we introduce TennisTV, the first high-quality benchmark specifically designed for understanding tennis rally videos. TennisTV defines a stroke event as the atomic unit, spanning movement initiation, stroke action choice, directional execution, and the ball’s landing. Each rally video is modeled as a time-ordered rally sequence composed of these consecutive stroke events. A total of 9 evaluation tasks that match the sports characteristics of tennis, including two types of multiple-choice questions and judgment statements. We design an automated annotation pipeline. Through three-stage annotation process, we realize structured annotation integration for the entire rally based on the rough annotation provided by the original data source[5], extract additional athletes’ movement information, and generate questions. Finally, through manual quality control, we construct the TennisTV benchmark, which contains 2943 curated questions.
@@ 2.1 Sports Understanding
Sports understanding [6] is a rapidly evolving field that encompasses multiple research topics and integrates diverse modalities, covering tasks such as action recognition [7], athlete analysis [8], tactical planning [9], foul recognition [10]. Recent work on MLLMs [11, 12, 13] has shifted toward general frameworks spanning diverse sports tasks, while existing work[5] in tennis primarily uses conventional neural networks for stroke label prediction, with limited modeling of rally level context and causal relations. Consequently, MLLM capabilities in fast-paced, high-frequency sports video remain largely unexplored. This paper uses tennis as the empirical domain and proposes a novel benchmark for fast-paced, high-frequency sports video understanding, with the goal of advancing future research in this area.
@@ 3 TennisTV Benchmark
To rigorously evaluate the performance of MLLMs in tennis video understanding, we introduce Tennis Tour Video (TennisTV). The benchmark comprises 9 subtasks with 2943 questions, supporting multiple levels of video understanding. As shown in Fig 2, this section outlines benchmark overview introduction and the automatic annotation pipeline.
Table 1: Main results on TennisTV. AvgSE and AvgRS denote the average metrics of tasks related to stroke event and rally sequence.
Method Frame Sampling MG AR HD HO AvgSE TP RC TI PW ErrA AvgRS Avg
Random - 20.0 25.0 20.0 33.3 24.6 25.0 25.0 25.0 50.0 50.0 35.0 30.4
Human-level - 47.8 100.0 96.3 93.3 84.4 91.6 92.5 98.0 94.3 87.5 92.8 89.0
@@
Video-LLaVA-7B 8 Frames 19.4 25.2 16.3 32.3 23.3 24.8 23.3 25.0 50.0 50.0 34.6 29.6
LLaVA-OneVision-7B 64 Frames 14.4 36.3 30.0 41.0 30.4 35.7 25.3 41.0 49.0 50.0 40.2 35.9
mPLUG-Owl3 128 Frames 16.1 42.8 33.8 46.0 34.7 35.7 18.3 40.7 50.0 50.0 38.9 37.0
Qwen2VL-7B 32 Frames 19.4 29.5 34.5 40.3 30.9 43.5 24.8 44.7 50.0 51.9 43.0 37.6
InternVideo2.5-8B 128 Frames 20.0 49.5 17.8 43.0 32.6 42.5 19.0 43.3 50.3 49.3 40.9 37.2
InternVL3-8B 64 Frames 31.9 34.8 31.3 45.3 35.8 40.7 30.3 44.3 50.0 50.0 43.1 39.8
Qwen2.5VL-3B 32 Frames 17.8 39.7 32.0 40.3 32.5 39.8 20.5 46.3 50.0 50.0 41.3 37.4
Qwen2.5VL-7B 32 Frames 26.7 48.6 34.5 45.7 38.9 42.2 22.3 42.7 48.0 51.0 41.2 40.2
MiMo-VL-7B SFT-2508 32 Frames 22.8 45.5 31.8 38.3 34.6 44.7 31.8 41.0 46.0 52.6 43.2 39.4
@@
Video-R1 32 Frames 21.7 45.9 28.8 42.0 34.6 39.4 28.3 40.7 52.0 47.4 41.6 38.5
VCR-Thinking-7B 32 Frames 28.3 45.5 35.5 23.7 33.3 38.5 33.3 44.0 47.7 50.0 42.7 38.5
VideoRFT 32 Frames 28.3 46.8 35.3 45.0 38.9 37.3 25.5 45.3 45.7 49.8 40.7 39.9
VersaVid-R1 32 Frames 24.4 52.3 36.8 44.7 39.6 37.9 27.0 46.7 50.0 51.2 42.6 41.2
MiMo-VL-7B RL-2508 32 Frames 22.8 48.0 33.3 42.7 36.7 36.7 28.8 49.3 57.3 49.5 45.3 41.5
@@ 3.1 Benchmark Task Analysis
Stroke Event. The stroke level perspective targets fine-grained phenomena, where the model must recognize and localize specific stroke events such as Action Recognition (AR), Movement Grounding (MG), Hit Direction (HD) and Hit Outcome (HO). These tasks compel the model to detect subtle spatiotemporal cues at the frame level during rapid, interactive exchanges.
Rally Sequence. The rally level perspective centers on the macro dynamics of a point and encompasses Technique Identification (TI), Rally Count (RC), Technique Preference (TP), Point Winner determination (PW) and Error Analysis (EA). These tasks require a model to analyze the entire game process, so as to grasp the overall trend of the rally and the players’ technical tendencies.
Overall, TennisTV provides a comprehensive and challenging benchmark. From a general understanding perspective, it encompasses both the perception of basic object motion information (HD, MG) and the comprehensive summarization of video events (RC). Within the sports domain, it further includes tactical understanding (TP) and the win or loss judgment based on tennis rules (PW).
@@ 3.2 Dataset Annotation Pipeline
TennisTV is derived from F3Set[5], which contains manually annotated shots from tennis tournament videos. Based on raw annotations, we perform a three-stage data annotation pipeline.
Basic Data Preparation We first aggregated the shot annotations corresponding to the same rally in temporal order and generated basic rally level technical statistics, such as rally length and the distribution of different strokes used by both players, to construct structured rally information. Next, we removed videos with an insufficient number of rallies from the original dataset. Moreover, because tennis tournaments are played on courts with diverse surfaces, the varying background colors can reduce viewing clarity. To address this issue, we employed Qwen2.5VL-7B[20] as a video quality judge model to filter out videos with poor visual quality.
@@
Question Generation We employed GPT-4o to aggregate the structured rally information and generated a comprehensive rally sequence description for each video. Using the rally descriptions and basic statistics, we generated 9 task questions using either GPT-4o or a predefined template. All questions were manually verified, resulting in a total of 2943 questions across 1392 videos.
@@ 4.1 Implementation Details
We evaluate 14 open-source MLLMs that can process multi-frame video inputs, including 9 non-thinking models: Video-LLaVA [17] , LLaVA-OneVision-7B, Qwen2.5VL-3B&7B [20] , Qwen2VL-7B, mPLUG-Owl3 [21] , InternVideo2.5 [2] and MiMoVL-SFT-7B-2508 [22] , as well as 5 thinking models: Video-R1 [23] , VideoChat-R1 [3] , VideoRFT [4] , VersaVid-R1 [19] and MiMoVL-RL-7B-2508 [22] . In addition, we assess GPT-4o, GPT-4.1 and Gemini2.5-Pro through their APIs. For the open-source models, we followed the input frame settings in VideoMME [24] . For the closed-source models, we adopted a video sampling rate of 1 FPS for evaluation due to budget consideration.
@@ 4.2 Evaluation Metric
We adopt two different metrics to evaluate the model performance on TennisTV. For PW and EA, the model is provided with a statement describing the outcome of the rally. The model’s answer is considered correct only if it correctly classifies the statement as True or False. For other tasks, we formulate them as multiple-choice questions and evaluate whether the model’s selected option is correct.
@@ 4.3 Main Results
Results Table 3 reports results on TennisTV. Among closed-source models, GPT-4.1 rank first with an overall score of 46.2%, narrowly surpassing GPT-4o and Gemini 2.5 Pro. Open-source systems also demonstrate strong competitiveness: under the non-thinking setting, Qwen2.5-VL-7B achieved a score of 40.2%, while the best thinking model MiMo-VL-7B improved this to 41.5% (+1.3pp). On average, open-source reasoning models outperform their non-reasoning counterparts but still lagged considerably behind leading closed-source commercial systems. Despite these gains, overall accuracy remains well below human performance and, on some tasks, even falls beneath the random-chance baseline, indicating substantial headroom for progress.
Fig. 3: Correlation matrix between 9 tasks in TennisTV.
@@
1 FPS 37.9 41.3 39.8
16 Frames 38.0 41.7 40.0
32 Frames 38.9 41.2 40.2
64 Frames 39.1 39.4 39.3
Finding II — The optimal frame sampling density varies across perceptual levels. As we discussed above, In fast-paced, high-frequency sports videos, the number of frames required varies across different levels of understanding. As shown in Table 2, performance with Qwen2.5-VL-7B peaks at 32 frames, while adding more frames yields diminishing or even negative returns, particularly for rally-sequence metrics. These results suggest that finer temporal resolution is most beneficial for capturing brief, high-frequency stroke cues, whereas rally level reasoning saturates earlier and is more sensitive to redundancy.
Fig. 4: Visualization of performance on three TennisTV tasks, which require no prior knowledge of tennis, based on the same base model.
Finding III — Reasoning not always helps. To disentangle the effect of reasoning from tennis-specific priors, we analyze three knowledge-light tasks (MG, HD, RC) that minimize dependence on specialized rules. We analyze Qwen2.5-VL-7B, the zero-shot chain-of-thought (CoT) prompting method [25] and three RL post-trained reasoning variants (VideoChat-R1, VideoRFT, VersaVid-R1). As shown in Fig. 4, reasoning variants exhibit consistent gains over the non-reasoning baseline on RC and HD, where multi-step aggregation and explicit trajectory reasoning are critical. However, on MG task, which focuses on temporal grounding, reasoning model do not show consistent improvement over the baseline. This suggests that existing RL strategies primarily enhance general content understanding, while temporal perception remains a key shortcoming.
@@ 5 Takeaways
  • Enhancing the temporal understanding ability of video reasoning models. The experimental results have proved that reasoning can improve rally level understanding, but it has not benefited from fine-grained temporal grounding task. This suggests that future RL-based post-training should emphasize temporal understanding and alignment mechanisms rather than relying solely on general video reasoning.
@@ 6 Conclusion
In this paper, we present TennisTV, the first benchmark for tennis video understanding tailored to evaluate MLLMs. TennisTV models each rally as a time-ordered sequence of consecutive stroke events, using automated pipelines for filtering and question generation. Unified evaluations over diverse MLLMs show that reasoning and careful frame sampling help, yet models still struggle with the sport’s fast, information-dense dynamics. We expect TennisTV to spur advances in hierarchical temporal reasoning and robust sports video understanding.